Consider a market with the demand curve given by ?=400−2?.
a. Suppose this market is monopolistic with a monopoly having a total cost function given by ??=1200+80?. Compute the market price and quantity sold in this market, as well as the level of profits for the monopolist.
b. Suppose this market is a perfect competition with a typical firm having a total cost function given by ??=1200+80?. Compute the market price and quantity sold in this market, as well as the level of profits (losses) for the firm.
c. Based on your findings in part (b), determine the behavior of the firm in the long-run.
Answer #1

A)

Q = 400-2P

Or, 2P = 400-Q

Or, P = 200-0.5Q

MR = dPQ/dQ = 200-Q

TC = 1200+80Q

MC = dTC/dQ = 80

A monopolist sets prices at the point MR = MC

That is,

200-Q = 80

Q1 = 120

P1 = $140

Profit = PQ - TC = (120*140) - (1200+80*120) = $6000

B)

A perfectly competitive firm sets prices at the point P = MC

That is,

P = 200-0.5Q = 80

Q2 = 240

P2 = 80

Profit = PQ - TC = (240*80) - (1200+80*240) = -$1200

C)

Since profits are negative, firms will start exiting the market in the long run, till the point profits fall to normal level.